Please show your work with the formulas written out. Please answer as if you can only use a four function calculator (because that is how I have to learn it). Do not just put down the calculator keystrokes I need to see every step and number to learn how to do it.

13. If you want to accumulate $250 after 30 years and your bank pays 2%, how much will you have to deposit monthly?

14. If you borrow $250 for 25 years at 6%, what will be your monthly payment?

15. If you can earn 15%, how long will it take for an investment today of $700 to grow to $5000?

16. If you save $100 monthly and your bank promises 15%, how many months will it take to accumulate $25,000?
Answer #1

13. FV =250
Number of Periods =30*12 =360
Rate per month =2%/12
Monthly Payment =FV/((1+r)^n-1)/r) =250/(((1+2%/12)^360-1)/(2%/12))=0.51

14. PV =250
Number of Periods =25*12 =300
Rate per month =6%/12 =0.5%
Monthly Payment =PV/(((1-(1+r)^-n)/r)=250/((1-(1+0.5%)^-300)/0.5%) =1.61